# A Prospect-Theoretic Policy Gradient Algorithm for Behavioral Alignment in Reinforcement Learning

### v2

@@ H.2 Alternative Practical Procedure for Computing Stochastic Policy Gradients
[…]
@@
Let $w$ be piecewise quadratic: there exists $q_{1}<q_{2}<....<q_{k}$, with $q_{1}=0$ and $q_{k}=1$, as well as reals $\alpha_{1},....,\alpha_{k}$, $\beta_{1},....,\beta_{k}$ and $\delta_{1},....,\delta_{k}$ such as $w(x)=\sum_{i=1}^{k-1}\mathds{1}_{[q_{i},q_{i+1}[}(t)(\frac{1}{2}\alpha_{i}t^{2}+\beta_{i}t+\delta_{i})$. For all $11\leq i\leq k-$ 1, define the $i$-th quantile of $\mathcal{U}(X)$ as $\tilde{q}_{i}:=\sup\{t\in\mathbb{R}\cup\{+\infty,-\infty\},\mathbb{P}(\mathcal{U}(X)>t)\geq q_{i}\}$. Then, for any given $t\in[\tilde{q}_{j+1},\tilde{q}_{j}[$:
@@
[…]
@@
Suppose that the weight function $w:[0,1]\mapsto[0,1]$ is piecewise affine, i.e. there exists $q_{1}<q_{2}<....<q_{k}$, with $q_{1}=0$ and $q_{k}=1$, as well as reals $\beta_{1},....,\beta_{k}$ and $\delta_{1},....,\delta_{k}$ s.t. $w(x)=\sum_{i=1}^{k-1}\mathds{1}_{[q_{i},q_{i+1}[}(x)(\beta_{i}x+\delta_{i})$ for any $x\in[0,1]$. Let $\tilde{q}_{i}:=\sup\{t\in\mathbb{R}\cup\{+\infty,-\infty\},\mathbb{P}(\mathcal{U}(X)>t)\geq q_{i}\}$ for any $i=1,\cdots,k\,.$ Then for any $11\leq j\leq k-$ 1 and any $t\in[\tilde{q}_{j+1},\tilde{q}_{j}[$,
@@
Appendix I More Details about Section 5 and Additional Experiments
Environment Utility function $w^{+}$ function Figure Comment
Grid Various 3-segment piecewise affine function Fig. 2 and Apdx I.4 We observe convergence and various behaviours for various utility functions
Traffic Control Identity Risk averse ($w_{ra}$) Apdx. I.5 The policy goes around the city center
Traffic Control Identity Risk neutral (Id.) Apdx. I.5 The policy goes through the city center
Traffic Control Identity Risk averse ($w_{ra}$) / Risk-neutral (Id.) Apdx. I.5 Same behavior, entropy regularization needed
Scalable Grid Identity Risk-averse ($w_{ra}$) Fig. 3 Our algorithm converges faster than CPT-SPSA-G for large grids
Electricity Management Identity Very risk averse ($w_{vra}$) / Very risk seeking ($w_{vrs}$) / Risk-neutral (Id.) Fig. 2 Convergence to different reward distributions in accordance to behavior to risk
Fig. 16 Exponential, Kahneman-Tversky Risk-neutral (Id.) Fig. 21 The result illustrates Theorem 4
Fig. 14(a) Identity Risk-seeking ($w_{rs}$) Fig. 15 The result illustrates Proposition 2
Table 2: Summary of experiments
@@ Appendix I More Details about Section 5 and Additional Experiments
Table 2 recaps the various experimental settings. The risk-neutral $w^{+}$ function is simply the identity function. As for the definition of other probability distortion functions $w^{+}$ we use for experiments, we define:
@@
Instead of vanilla stochastic gradient descent, we use the Adam optimizer to speed up convergence. In our Python implementation, we use the same batch of trajectories for estimating the function $\varphi$ and for the performing the stochastic gradient ascent step. We have run the experiments on a laptop with a 13th Gen Intel Core i7-1360P2.20 GHz CPU and 32 GB of RAM.
We use the tanh activation function before the last softmax layer to encourage exploration and reduce the risk of converging to local optima which may occasionally occur for some runs.
I.1 Additional Figure
@@
I.2 Illustration of Proposition 2: about the need for Stochastic Policies in CPT-RL
@@
We illustrate Proposition 2 and study experimentally the behavior of our algorithm with regards to small batch sizes.
Setting. We use the barebones setting (Figure 14) introduced in the proof of Proposition 2 with its $w$ function that aggressively focuses on the 10% of favorable outcomes. Denoting by $p$ the probability of choosing A for a given policy, we look at the value of $p$ at convergence (1000 optimization steps) for various batch sizes. The optimal policy corresponds to $p=0.8$.
Figure 15: Results of the experiment on non-deterministic policies and batch size influence, over $100$ runs. The black dots are the medians and the shaded area represents the interquartile range.
Insights. For each batch size we test, we run and plot a hundred training rounds (Figure 15). We fist observe that the policy we obtain with our algorithm indeed approaches the optimal $p=0.8$ policy. The estimation error (w.r.t. the optimal theoretical value of $p=0.8$) appears to be of order 1batch $\frac{1}{\sqrt{\text{batch size}}}$. It was to be expected that a small batch size would lead to a bias in CPT value and CPT gradient estimation, and, finally, in policy, as a small batch size renders impossible an accurate estimation of the probability distribution of the total return function. The fact that this bias appears to be proportional to the inverse of the square root of the batch size is in line with the standard statistical intuition (as e.g. per the central limit theorem). In our particular example, the estimated $p$ is below (and not above) the theoretical $p$. This is likely because our $w$ function places a strong weight on the top $10\%$ of outcomes. Hence there is an imbalance between the impact of overestimating and underestimating the proportion of good outcomes in a given run: if we underestimate the probability of getting $+1.5$ with a given policy due to sampling, the effect will be stronger than the opposite effect we would get by overestimating the probability of the same error. As the batch size grows, the estimation error is reduced and the effect vanishes.
I.3 Illustration of Theorem 4: Markovian vs Non-Markovian Policies for CPT-RL
@@
To illustrate the fundamental difference between memoryless utility functions studied in Theorem 4 and the others we conduct a small experiment on a simple setting (Figure 16), similar to the one introduced in the proof of the theorem. We consider three states and three actions. From the starting state, any action leads to the second state with probability $1$ and yields a reward of $+1$ with probability $\frac{1}{2}$ and of $-1$ with probability $\frac{1}{2}$. Once in the second state, the first action yields reward $+1$ with probability $1$, the second action yields 00 or $3$ with probability $\frac{1}{2}$ each, and the third action always yields 00. We compare the performance of a policy parametrized in $\Pi_{\Sigma,NS}$ and one in $\Pi_{M,NS}$.
@@
Figure 17: Comparison of Markovian and Non-Markovian policy performances for non-exponential (left) and exponential (right) utility functions. Shaded areas represent a range of $\pm$ one standard deviation over $20$ runs.
I.4 Grid Environment
@@
Influence of the utility function. We consider a 4x4 grid for our illustration purpose. Our agent starts on a random square on one of the three upper rows of the grid, and can move in all four directions. Any move to an empty square will award it a random reward of $-1$ with probability $\frac{1}{2}$ and of $+0.8$ with probability $\frac{1}{2}$. Therefore, longer trajectories are slightly costly in expectation, and generate significant variance. In two corners of the grid, we add cells that yields rewards of $+5$ for one or $+6$ for the other, and conclude the episode. Illegal moves (attempting to leave the grid) are punished by a negative reward. Our parameterized policy is a neural network whose last layer is activated with softmax and has 4 coordinates corresponding to the 4 different possible moving actions. We consider solving (CPT-PO) with different utility functions: risk-neutral identity utility, risk-averse KT utility, as well as exponential utility function. The obtained policies differ depending on the utility function. For examples of risk-neutral/averse policies obtained, see Fig. 19(b).
I.5 Traffic Control
@@
[…]

### v3

@@ I.2 Alternative Practical Procedure for Computing Policy Gradients
$\int_{a}^{b}w^{\prime}(\mathbb{P}(\mathcal{U}(X)>z)dz$ $=\int_{0}^{b-a}w^{\prime}(\mathbb{P}(\mathcal{U}(X)-a>v))dv$
$=\int_{0}^{b-a}(\alpha(\mathbb{P}(\mathcal{U}(X)-a>v))+\beta)dv$
$=\alpha\int_{0}^{b-a}\mathbb{P}(\mathcal{U}(X)-a>v)dv+\beta(b-a)$
$=\alpha\int_{0}^{b-a}\mathbb{P}(Y_{a,b}>v)dv+\beta(b-a)$
$=\alpha\int_{0}^{+\infty}\mathbb{P}(Y_{a,b}>v)dv+\beta(b-a)$
@@
Let $w$ be piecewise quadratic: there exists $q_{1}<q_{2}<....<q_{k}$, with $q_{1}=0$ and $q_{k}=1$, as well as reals $\alpha_{1},....,\alpha_{k}$, $\beta_{1},....,\beta_{k}$ and $\delta_{1},....,\delta_{k}$ such as $w(x)=\sum_{i=1}^{k-1}\mathds{1}_{[q_{i},q_{i+1}[}(t)(\frac{1}{2}\alpha_{i}t^{2}+\beta_{i}t+\delta_{i})$. For all $1\leq i\leq k-1$, define the $i$-th quantile of $\mathcal{U}(X)$ as $\tilde{q}_{i}:=\sup\{t\in\mathbb{R}\cup\{+\infty,-\infty\},\mathbb{P}(\mathcal{U}(X)>t)\geq q_{i}\}$. Then, for any given $t\in[\tilde{q}_{j+1},\tilde{q}_{j}[$:
@@
$\int_{0}^{t}w^{\prime}(\mathbb{P}(U(X)>z)dz$ $=\sum_{i=j+1}^{k-1}\int_{\tilde{q}_{i+1}}^{\tilde{q}_{i}}w^{\prime}(\mathbb{P}(U(X)>z))dz+\int_{\tilde{q}_{j+1}}^{t}w^{\prime}(\mathbb{P}(U(X)>z))dz$
@@
Suppose that the weight function $w:[0,1]\mapsto[0,1]$ is piecewise affine, i.e. there exists $q_{1}<q_{2}<....<q_{k}$, with $q_{1}=0$ and $q_{k}=1$, as well as reals $\beta_{1},....,\beta_{k}$ and $\delta_{1},....,\delta_{k}$ s.t. $w(x)=\sum_{i=1}^{k-1}\mathds{1}_{[q_{i},q_{i+1}[}(x)(\beta_{i}x+\delta_{i})$ for any $x\in[0,1]$. Let $\tilde{q}_{i}:=\sup\{t\in\mathbb{R}\cup\{+\infty,-\infty\},\mathbb{P}(\mathcal{U}(X)>t)\geq q_{i}\}$ for any $i=1,\cdots,k\,.$ Then for any $1\leq j\leq k-1$ and any $t\in[\tilde{q}_{j+1},\tilde{q}_{j}[$,
@@
Appendix J Proofs for Section 4
The proofs in this section use similar techniques as the proofs developed in L.A. et al. [2016]. Nevertheless, our policy gradient estimator is different and based on first-order information (using our policy gradient theorem) and this entails a few technical differences in the analysis.
J.1 Proof of Proposition 5: Consistency of the PG estimator
The proof of this result follows similar lines to the proof of Proposition 3 in L.A. et al. [2016] with the difference that we have a difference policy gradient estimator. The proof is based on applying a dominated convergence theorem (Theorem 17 below).
Observe first using the definition of $\hat{\phi}_{n}^{+}$ (see Algorithm 1) that we can write:
[…]
$=\int_{0}^{R(\tau)}w^{\prime}_{+}(1-\hat{F}_{n}^{+}(x))dx\,$, (15)
where $\hat{F}_{n}^{+}(x)$ is the empirical distribution of $u^{+}(R(\tau))$ (where $\tau$ is a random trajectory).
Using this form, it remains to prove that $\int_{0}^{R(\tau)}w^{\prime}_{+}(1-\hat{F}_{n}^{+}(x))dx\to\int_{0}^{R(\tau)}w^{\prime}_{+}(\mathbb{P}(u^{+}(R(\tau)>x))dx$ using Theorem 17 with $f_{n}=w_{+}^{\prime}(1-\hat{F}_{n}^{+}(x))$ and $g_{n}=L_{w}(1-\hat{F}_{n}^{+}(x))$ and using the fact that $g_{n}$ converges almost surely to $\mathbb{P}(u^{+}(R(\tau)>x)$ uniformly in $x$ by the Glivenko-Cantelli theorem. See proof of Proposition 3 in L.A. et al. [2016] p. 20 for details.
Theorem 17.
(Generalized Dominated Convergence theorem) Let $(f_{n})$ be a sequence of measurable functions on $E$ that converge pointwise a.e. on a measurable space $E$ to $f$. Suppose there exists a sequence $(g_{n})$ of integrable functions on $E$ that converge pointwise almost everywhere on $E$ to $g$ such that $|f_{n}|\leq g_{n}$ for all $n\in\mathbb{N}$. If $\lim\limits_{n\rightarrow\infty}$ $\int_{E}$ $g_{n}$ = $\int_{E}$ $g$, then $\lim\limits_{n\rightarrow\infty}$ $\int_{E}$ $f_{n}$ = $\int_{E}$ $f$.
J.2 Proof of Proposition 6: Sample complexity of the PG estimator
Observe first that our policy gradient estimator can be written as follow:
$\hat{\nabla}_{n,m}J(\theta)$ $=Y_{n}\cdot\left(\frac{1}{m}\sum_{l=1}^{m}X_{l}\right)\,,\quad\text{where}\quad Y_{n}:=\hat{\phi}_{n}^{+}-\hat{\phi}_{n}^{-}\,,\quad X_{l}:=\sum_{t=0}^{H-1}\nabla_{\theta}\log\pi_{\theta_{k}}(a_{t}^{l}|h_{t}^{l})\,$, (16)
see Algorithm 1. The idea of the proof is to show that both estimators in the policy gradient estimator concentrate around their mean. On the one hand, for the random variable $Y_{n}$ we will use the following Dvoretzky-Kiefer-Wolfowitz inequality (see Lemma 18). On the other hand, for the variable $X_{l}$ which is vector-valued, we will use a concentration inequality for bounded vector-valued random variables from the literature [Jin et al., 2019].
We start with the random variable $Y_{n}$ and show how to control $\hat{\phi}_{n}^{+}$, $\hat{\phi}_{n}^{-}$ can be bounded in the same way. Recalling the rewritten form of $\hat{\phi}_{n}^{+}$ in (J.1), it appears that $\hat{\phi}_{n}^{+}$ is clearly comparable to $\varphi^{+}(v):=\int_{z=0}^{\max(v,0)}w_{+}^{\prime}(\mathbb{P}(u^{+}(R(\tau))>z))dz$ as defined in Theorem 3.
The rest of the proof for controlling this quantity follows similar lines to the proof of Proposition 3 in L.A. et al. [2016]. Since $u^{+}(\tau)$ is supposed to be bounded by $M_{u}$ and $w_{+}^{\prime}$ (note here this is the derivative of $w_{+}$ unlike in [L.A. et al., 2016, Proposition 3]) is $L_{w}$-Lipschitz, we have
$|\hat{\phi}_{n}^{+}-\varphi^{+}(R(\tau))|$ $=\left|\int_{0}^{R(\tau)}w_{+}^{\prime}(\mathbb{P}(u^{+}(R(\tau)))>z)dz-\int_{0}^{R(\tau)}w_{+}^{\prime}(1-\hat{F}_{n}^{+}(z))dz\right|$
$\leq\int_{0}^{M_{u}}\left|w_{+}^{\prime}(\mathbb{P}(u^{+}(R(\tau)))>z)-w_{+}^{\prime}(1-\hat{F}_{n}^{+}(z))\right|dz$
$\leq\int_{0}^{M_{u}}L_{w}\left|\mathbb{P}(u^{+}(R(\tau)))<z)-\hat{F}_{n}^{+}(z)\right|dz$
$\leq L_{w}M_{u}\sup_{z\in\mathbb{R}}\left|\mathbb{P}(u^{+}(R(\tau))<z)-\hat{F}_{n}^{+}(z)\right|.$
Using the DKW inequality (Lemma 18 below)yields:
$\mathbb{P}\left(|\hat{\phi}_{n}^{+}-\varphi^{+}(R(\tau))|>\epsilon/2\right)\leq\mathbb{P}\left(L_{w}M_{u}\sup_{z\in\mathbb{R}}\left|\mathbb{P}(u^{+}(R(\tau))<z)-\hat{F}_{n}^{+}(z)\right|>\epsilon/2\right)\leq 2e^{-n\frac{\epsilon^{2}}{2L_{w}^{2}M_{u}^{2}}}.$ (17)
A similar inequality holds for $|\hat{\phi}_{n}^{-}-\varphi^{-}(R(\tau))|$.
As for the random variable $X_{l}$, we observe that it is bounded by $HM_{g}$ under our score function boundedness assumption. Applying Corollary 7 in Jin et al. [2019], there exists an absolute constant $c>0$ such that with probability at least $1-\delta$ (for any $\delta\in(0,1)$), we have:
$\left\|\frac{1}{m}\sum_{l=1}^{m}X_{l}-\mathbb{E}[X_{1}]\right\|\leq c\sqrt{\frac{(HM_{\psi})^{2}\log\left(\frac{2d}{\delta}\right)}{m}}\,.$ (18)
Combining the above inequality with (17), we obtain with probability at least $1-\delta$,
$\|\hat{\nabla}_{n,m}J(\theta)-\nabla J(\theta)\|\leq cHM_{\psi}\sqrt{\frac{4L_{w}^{2}M_{u}^{2}\log\left(\frac{1}{\delta}\right)}{n}}+cM_{u}L_{w}\sqrt{\frac{(HM_{\psi})^{2}\log\left(\frac{2d}{\delta}\right)}{m}}\,.$ (19)
It remains to pick $n\geq\frac{(cHM_{\psi}M_{u}L_{w})^{2}\ln(1/\delta)}{\varepsilon^{2}}$ and $m\geq\frac{(cHM_{\psi}M_{u}L_{w})^{2}\ln(2d/\delta)}{\varepsilon^{2}}$ to conclude the proof.
Lemma 18.
(Dvoretzky-Kiefer-Wolfowitz (DKW) inequality) Let ${\hat{F}_{n}}(u)=\frac{1}{n}\sum_{i=1}^{n}1_{((u(X_{i}))\leq u)}$ denote the empirical distribution of a random variable $U$, with $u(X_{1}),\ldots,u(X_{n})$ being sampled from the random variable $u(X)$. Then, for any $n$ and any $\epsilon>0$, we have
$\mathbb{P}\left(\sup_{x\in\mathbb{R}}|\hat{F_{n}}(x)-F(x)|>\epsilon\right)\leq 2e^{-2n\epsilon^{2}}.$ (20)
J.3 Proof of Proposition 7: Asymptotic convergence
This result follows from a standard application of the stochastic approximation framework (see e.g. Borkar [2008], Benaïm [2006]) similarly to L.A. et al. [2016]. Under our regularity assumptions on the utility and weight functions, we can show asymptotic convergence of the iterates to the set of stationary points of our CPT objective. The idea is to show that the iterates track a gradient flow defined by the gradient of the CPT-PO objective with vanishing step sizes. The proof follows the same steps as the proof of Theorem 1 in L.A. et al. [2016] (see appendix D, p. 25), so we do not reproduce it here, up to the difference that our policy gradients are not estimated using zeroth-order information but rather using our PG Thm. 3 and Prop. 4.
To apply existing stochastic approximation results, it suffices to write:
$\theta_{k+1}=\theta_{k}+\alpha_{k}(\nabla J(\theta_{k})+\beta_{t}+\eta_{t})\,$, (21)
where $\eta_{t}$ is a martingale difference sequence, $\beta_{t}$ is a possible non-zero bias (in our case we use unbiased estimates) and $\alpha_{k}$ satisfies the Robbins-Monro conditions.
Appendix K More Details about Section 5 and Additional Experiments
Batch size and quantile estimation. Quantile estimation is generally more sensitive to batch size than expectation-based methods. However, in practice, we observed that relatively small batches (5–32 samples) performed robustly across all environments. As shown in Fig. 15, larger batches can improve performance, but small batches already yield competitive CPT values.
History dependence of policy networks. In most experiments, we used feed-forward networks without explicit history dependence. We found that Markovian policies performed well in our simple simulations for applications like financial markets and Mujoco control. In other domains (e.g., electricity management), we incorporated partial history by expanding the input window to represent multiple past states. Typically, the input size was chosen to capture sufficient context without encoding the full trajectory length ($H$). This limited form of history dependence was sufficient in our settings, though incorporating more expressive temporal models could enhance performance further.
Hardware and execution time. We have run part of the experiments on a laptop with a 13th Gen Intel Core i7-1360P2.20 GHz CPU and 32 GB of RAM, and the remaining simulations (section 5, App. K.8, K.9) on a MacBook Pro M4 with 48 GB of RAM. Experiments can take a few seconds for the simplest ones, to a few minutes for the ones in the main part. More complex ones take a couple of hours, e.g. for electricity management (App. K.7), Finance (App. K.8) and MuJoCo (App. K.9).
Software. Our experiments are coded in Python and mainly use Pytorch [Paszke et al., 2019].
@@
Weight functions. The risk-neutral $w^{+}$ function is simply the identity function. As for the definition of other probability distortion functions $w^{+}$ we use for experiments, we define:
@@
Optimizer. Instead of vanilla stochastic gradient descent, we use the Adam optimizer to speed up convergence. In our Python implementation, we use the same batch of trajectories for estimating the function $\varphi$ and for performing the stochastic gradient ascent step.
Neural net activation function. We use the tanh activation function before the last softmax layer to encourage exploration and reduce the risk of converging to local optima which may occasionally occur for some runs.
K.1 More details about bandit simulations in Section 5
In addition to the parameters specified in the main part, we used the following set of parameters to conduct the bandit experiments in Section 5:
Table 2: Hyperparameters for the bandit experiments in Section 5
Hyperparameter Value
Optimizer Adam
Learning rate 0.01
Number of episodes 1000
Batch size 10
We used a piecewise linear probability weight function given by the coefficients: $w=[1.0,0.0,1.0,0.0,0.5]$ where the notation $a_{1},a_{2},\dots,a_{i},b_{1},b_{2},\dots,b_{i},c_{1},\dots,c_{i-1}])$ refers to a piecewise affine $w$ function with $f(x)=a_{i}x+b_{i}$ for $c_{i-1}<x<c_{i}$.
In Fig. 12, we show a case where CPT-PG learns a stochastic policy.
Figure 12: Comparison of our CPT-PG algorithm with exponential risk-sensitive PG and vanilla PG on a simple 2-action bandit setting. (Lower figure) Loss lottery setting: Only CPT picks the risky action with high probability. Note here that the policy trained by CPT-PG is stochastic.
K.2 CPT compared to Distortion Risk Measures
@@
K.3 Illustration of the need for stochastic policies in CPT-RL
@@
We study experimentally the behavior of our algorithm with regards to small batch sizes.
Setting. We use the barebones setting (Figure 14) with a $w$ function that aggressively focuses on the 10% of favorable outcomes. Denoting by $p$ the probability of choosing A for a given policy, we look at the value of $p$ at convergence (1000 optimization steps) for various batch sizes. The optimal policy corresponds to $p=0.8$.
Figure 15: Results on non-deterministic policies and batch size influence, over $100$ runs. The black dots are the medians and the shaded area represents the interquantile range.
Insights. For each batch size we test, we run and plot a hundred training rounds (Figure 15). We fist observe that the policy we obtain with our algorithm indeed approaches the optimal $p=0.8$ policy. The estimation error (w.r.t. the optimal theoretical value of $p=0.8$) appears to be of order $\frac{1}{\sqrt{\text{batch size}}}$. It was to be expected that a small batch size would lead to a bias in CPT value and CPT gradient estimation, and, finally, in policy, as a small batch size renders impossible an accurate estimation of the probability distribution of the total return function. The fact that this bias appears to be proportional to the inverse of the square root of the batch size is in line with the standard statistical intuition (as e.g. per the central limit theorem). In our particular example, the estimated $p$ is below (and not above) the theoretical $p$. This is likely because our $w$ function places a strong weight on the top $10\%$ of outcomes. Hence there is an imbalance between the impact of overestimating and underestimating the proportion of good outcomes in a given run: if we underestimate the probability of getting +1.5+1.5 with a given policy due to sampling, the effect will be stronger than the opposite effect we would get by overestimating the probability of the same error. As the batch size grows, the estimation error is reduced and the effect vanishes.
K.4 Markovian vs Non-Markovian Policies for CPT-RL
@@
To illustrate the fundamental difference between memoryless utility functions and others we conduct a small experiment on a simple setting (Figure 16), similar to the one introduced in the proof of the theorem. We consider three states and three actions. From the starting state, any action leads to the second state with probability $1$ and yields a reward of +1+1 with probability $\frac{1}{2}$ and of 1-1 with probability $\frac{1}{2}$. Once in the second state, the first action yields reward +1+1 with probability $1$, the second action yields 0 or $3$ with probability $\frac{1}{2}$ each, and the third action always yields 0. We compare the performance of a policy parametrized in $\Pi_{\Sigma,NS}$ and one in $\Pi_{M,NS}$.
@@
Figure 17: Comparison of Markovian and Non-Markovian policy performances for non-exponential (left) and exponential (right) utility functions. Shaded areas represent a range of $\pm$ one standard deviation over 2020 runs.
K.5 Grid Environment
@@
Influence of the utility function. We consider a 4x4 grid for our illustration purpose. Our agent starts on a random square on one of the three upper rows of the grid, and can move in all four directions. Any move to an empty square will award it a random reward of 1-1 with probability $\frac{1}{2}$ and of +0.8+0.8 with probability $\frac{1}{2}$. Therefore, longer trajectories are slightly costly in expectation, and generate significant variance. In two corners of the grid, we add cells that yields rewards of +5+5 for one or +6+6 for the other, and conclude the episode. Illegal moves (attempting to leave the grid) are punished by a negative reward. Our parameterized policy is a neural network whose last layer is activated with softmax and has 4 coordinates corresponding to the 4 different possible moving actions. We consider solving CPT-PO with different utility functions: risk-neutral identity utility, risk-averse KT utility, as well as exponential utility function. The obtained policies differ depending on the utility function. For examples of risk-neutral/averse policies obtained, see Fig. 19(b).
K.6 Traffic Control over a Grid
We simulate a car agent navigating a city grid, where central roads are faster but risk higher delays (see fig. 20). The agent must balance speed against risk by avoiding the city center. In risk-neutral settings, the agent favors faster routes, while risk-averse policies avoid the risky central roads. Fig. 20 (center) demonstrates that our algorithm successfully adapts to different risk-weighted objectives. We also consider solving CPT-PO with different utility functions: risk-neutral identity utility, risk-averse KT utility, as well as exponential utility function. Fig. 23 show the corresponding different CPT returns observed. The obtained policies differ depending on the utility function. For examples of risk-neutral/averse policies obtained, see Fig. 19(b) in appendix K.5.
Figure 20: Traffic control environment: red roads in the city center are prone to congestion.
@@
0 $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$ $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$ $\downarrow$